How much momentum will a dumb-bell of mass 10 kg transfer to the floor if it falls from a height of 80 cm? Take its downward acceleration to be 10 ms–2,

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Mass of dumb-bell = 10 kg
Height, h = 80 cm
= 0.8 m
a = 10 m/s^2
u = 0
v^2– u^2= 2 as
v^2– (0)^2= 2 × 10 × 0.8
v^2= 16
v = 4 m/s
Momentum p = mv
= 10 × 4
= 40 kgm/s


Thanks